OpenStudy (anonymous):

What is the perimeter of a regular pentagon with side length 7 cm?

OpenStudy (angela210793):

5*7=35cm

OpenStudy (anonymous):

35 cm

OpenStudy (ksaimouli):

35cm2

OpenStudy (angela210793):

P=x+x+x+x+x=5x x=7 P=5*7=35cm

OpenStudy (angela210793):

not cm^2 just cm

OpenStudy (ksaimouli):

ok